# Atlas3D: Physically Constrained Self-Supporting Text-to-3D for Simulation and Fabrication

### v1

@@ Abstract
Existing diffusion-based text-to-3D generation methods primarily focus on producing visually realistic shapes and appearances, often neglecting the physical constraints necessary for downstream tasks. Generated models frequently fail to maintain balance when placed in physics-based simulations or 3D printed. This balance is crucial for satisfying user design intentions in interactive gaming, embodied AI, and robotics, where stable models are needed for reliable interaction. Additionally, stable models ensure that 3D-printed objects, such as figurines for home decoration, can stand on their own without requiring additional supports. To fill this gap, we introduce Atlas3D, an automatic and easy-to-implement method that enhances existing Score Distillation Sampling (SDS)-based text-to-3D tools. Atlas3D ensures the generation of self-supporting 3D models that adhere to physical laws of stability under gravity, contact, and friction. Our approach combines a novel differentiable simulation-based loss function with physically inspired regularization, serving as either a refinement or a post-processing module for existing frameworks. We verify Atlas3D’s efficacy through extensive generation tasks and validate the resulting 3D models in both simulated and real-world environments. Project page: https://yunuoch.github.io/Atlas3D/.
@@ 2 Related Work
Most existing 3D generative models focus only on geometry or appearance modeling, with physics priors being underexplored. Time-independent physical constraints, such as penetrations, can be directly defined by penalties [97; 17; 82; 40; 102]. For time-dependent physical qualities, such as stability and comfort, data-driven quality checkers trained with offline simulators can be applied [10; 48; 4]. Offline simulators can also be used as validators to augment the training dataset [71] and update design with reinforcement learning [85], and as dynamics generators for generated 3D assets [91; 21; 62]. Another direction is to utilize differentiable simulations, which have be widely used in tasks like robotic control [74; 19; 61; 34; 76] or inverse problems [33; 20; 103; 53; 32; 73; 7; 104]. They can also be applied in 3D content generation to define physics-based losses to aid per-instance generation [54; 92] or model training [47].
@@ 3.1 Score Distillation Sampling
SDS-based methods have proven effective in distilling 3D models from 2D images. They utilize a 3D representation such as implicit density field, implicit Signed Distance Field (SDF), or tetrahedral SDF [67], a differentiable renderer like NeRF [50], NeuS [86], or Nvdiffrast [26], and a pre-trained text-to-image model such as Stable Diffusion [64] serving as diffusion guidance. The generation process involves optimizing the parameters $\theta$ of the 3D representation, where the 3D shape is differentiably rendered to 2D images ${\mathbf{z}}=g(\theta)$ and compared against the real distribution from the diffusion model with text guidance $y$. The score is backpropagated to the parameter space via:
$\nabla\mathcal{L}_{\text{SDS}}({\mathbf{z}}=g(\theta))=\mathbb{E}_{t,\epsilon}\left[w(t)(\hat{\epsilon}_{\phi}({\mathbf{z}}_{t},y,t)-\epsilon)\frac{\partial{\mathbf{z}}}{\partial\theta}\right],$ (1)
where ${\mathbf{z}}_{t}$ is the noisy image at noise level $t$, $w(t)$ is a weighting function, and $\hat{\epsilon}_{\phi}$ is the predicted noise. We refer readers to [58; 35; 46] for more technical details on SDS optimization.
@@ 4.1 Physics Incorporation
As mentioned in §$\S$ 3.2, we predict the dynamic behavior of generated models by rigid body simulations. While various explicit or implicit representations of 3D shapes can be chosen in a generation network, we opt for triangular meshes in our framework as they facilitate frictional contact modeling and simplify kinematics computation. Given a triangle surface mesh representation ${\mathbf{X}}(\theta)$, where $\theta$ is the implicit parameter, we integrate ${\mathbf{X}}$ into a rigid body represented by the dynamic state $\Psi(t)=[{\mathbf{T}}(t),{\mathbf{R}}(t),{\mathbf{P}}(t),{\mathbf{L}}(t)]^{T}$, where the world-space location ${\mathbf{x}}$ of any point ${\mathbf{X}}$ on the body is ${\mathbf{x}}(t)={\mathbf{R}}(t){\mathbf{X}}+{\mathbf{T}}(t)$. Assuming the 3D model is initially placed upright on the ground with the bottom point touching the surface, we define standability as:
@@ 4.1.2 Stable Equilibrium
Inspired by the concept of a potential well [13], we augment our optimization objective with a robust and efficient stable equilibrium loss $\mathcal{L}_{\text{stable}}$. Specifically, for an object to be robustly standable, it needs to reside at a local minimum of potential energy—specifically, gravitational potential energy—so that if perturbed, gravity will act as a restoring force that returns the object to its original state. For a rigid body, gravitational potential energy is determined by the height of its center of mass. Thus, for any object in a stable equilibrium state, the center of mass would rise if it is slightly perturbed. This leads to our formulation of the stable equilibrium loss $\mathcal{L}_{\text{stable}}$. Let $\hat{{{\mathbf{z}}}}$ denote the unit vector in the vertical direction (opposite to gravity) and $H({{\mathbf{x}}})$ denote the distance of the point ${{\mathbf{x}}}$ to the ground, assuming the object’s pivot point is at $z=0$. The stable equilibrium loss is defined as:
@@ 5.1 Simulation Verification
Figure 6: Success rate of models standing under perturbation.
@@
Since we do not assume a specific baseline model in the first stage, we can vary the model used in the coarse stage to generate versatile, physically-aware 3D meshes based on different existing SDS generation models. We use MVDream [69] as another baseline model and compare the results side-by-side with ours in Fig. 5. Our method improves the geometry of the mesh and ensures stability in simulation. The quality of the texture and the main part of the shape is determined by underlying model used in the coarse stage, while we focus on improving the physical stability in the refinement stage. More qualitative comparisons with previous methods are provided in the Appendix.
@@
To mimic this uncertainty in our framework, we evaluate the stability of our generated models under a small initial rotation. With a given precision $\phi_{\max}$, we rotate the generated mesh at random angles $\phi_{y},\phi_{x}\in(-\phi_{\max},\phi_{\max})$ in both the $y$ and $x$ axes, respectively (with the $z$ axis being the up direction). The mesh is then placed close to the ground and tested in a simulation to see if it can still stand. We choose 13 different values of the maximum perturbation angle $\phi_{\max}$ and perform 100 random tests with each angle on 6 of our generated models. We report the success rate in Fig. 6. We define successful test as: after a sufficiently long time period, the maximum height of the model stays within 3% of the initial maximum height.
@@
Due to the presence of our stability loss, our generated models can withstand small initial perturbations, while the baseline models generated without physical constraints fail to stand when placed straight up (without rotation), let alone with perturbations.
@@
To evaluate the standability of the baseline method and our method, we run rigid body simulation in Warp with simulation end time $T=2.0$ at which almost all objects have reached steady state. We propose Time-Averaged Rotation Deviation Loss (TRD) defined as
@@
to assess the standability, as a representation of the average tilting of the upward direction (of the object) over time. We approximate the integral (Eq. 11) with discrete quadrature $\Delta t=0.02.$ Results are plotted in Fig. 8  [^1] Results are ranked by TRD scores of the baseline approach.. The mean TRD score is reduced by more than six times compared with the baseline method. As shown in Table. 1, we calculate the average CLIP score [63] of 107 generated shapes for both our proposed method and baseline. Furthermore, Elo (GPT-4o) [90] scores are presented. Both metrics illustrate that our method not only implements physics-based stability but also maintains the fidelity of the generated 3D shapes in terms of content alignment and overall shape quality. More details are provided in the Appendix.
@@ Appendix A Appendix
A.1 Implementation and training details
@@
A.2 Additional results
A.2.1 Qualitative Comparison with Magic3D baseline
@@
A.2.2 More comparison with MVDream baseline
@@
A.2.3 Quantitative with Magic3D baseline
The results of the quantitative experiments are shown in 1. For CLIP score calculation, we specifically employ openai/clip-vit-large-patch14-336 as the check point of CLIP model. We rendered images from various angles (0-360 degree, 3 degree as the interval). For Elo (GPT-4o) benchmark, we made use of the newly released GPT-4o model instead of the one in [90] due to capacity limitation. Additionally, we reduce the number of views to 6 because of token length limitation.
A.3 Real world experiments
A.3.1 Robot manipulation results [^2] We define as upward pose for the kangaroo figure as its tail touching the ground, as a real kangaroo does.
@@ A.3 Real world experiments
A.3.2 User study results

### v2

@@ Abstract
Existing diffusion-based text-to-3D generation methods primarily focus on producing visually realistic shapes and appearances, often neglecting the physical constraints necessary for downstream tasks. Generated models frequently fail to maintain balance when placed in physics-based simulations or 3D-printed. This balance is crucial for satisfying user design intentions in interactive gaming, embodied AI, and robotics, where stable models are needed for reliable interaction. Additionally, stable models ensure that 3D-printed objects, such as figurines for home decoration, can stand on their own without requiring additional support. To fill this gap, we introduce Atlas3D, an automatic and easy-to-implement method that enhances existing Score Distillation Sampling (SDS)-based text-to-3D tools. Atlas3D ensures the generation of self-supporting 3D models that adhere to physical laws of stability under gravity, contact, and friction. Our approach combines a novel differentiable simulation-based loss function with physically inspired regularization, serving as either a refinement or a post-processing module for existing frameworks. We verify Atlas3D’s efficacy through extensive generation tasks and validate the resulting 3D models in both simulated and real-world environments.
@@ 2 Related Work
Most existing 3D generative models focus only on geometry or appearance modeling, with physics priors being underexplored. Time-independent physical constraints, such as penetrations, can be directly defined by penalties [97; 17; 82; 40; 102]. For time-dependent physical qualities, such as stability and comfort, data-driven quality checkers trained with offline simulators can be applied [10; 48; 4]. Offline simulators can also be used as validators to augment the training dataset [71] and update the design with reinforcement learning [85], and as dynamics generators for generated 3D assets [91; 21; 62]. Another direction is to utilize differentiable simulations, which have be widely used in tasks like robotic control [74; 19; 61; 34; 76] or inverse problems [33; 20; 103; 53; 32; 73; 7; 104]. They can also be applied in 3D content generation to define physics-based losses to aid per-instance generation [54; 92] or model training [47].
@@ 3.1 Score Distillation Sampling
SDS-based methods are shown to be effective in distilling 3D models from 2D images. They utilize a 3D representation such as implicit density field, implicit Signed Distance Field (SDF), or tetrahedral SDF [67], a differentiable renderer like NeRF [50], NeuS [86], or Nvdiffrast [26], and a pre-trained text-to-image model such as Stable Diffusion [64] serving as diffusion guidance. The generation process involves optimizing the parameters $\theta$ of the underlying 3D representation, where the 3D shape is differentiably rendered to 2D images $\bm{z}=g(\theta)$ and compared against the real distribution from the diffusion model with text guidance $y$:
$\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,\epsilon}\left[w(t)\left\|\hat{\epsilon}_{\phi}(\bm{z}_{t},y,t)-\epsilon\right\|^{2}\right],$ (1)
where $\bm{z}_{t}$ is the noisy image at noise level $t$, $w(t)$ is a weighting function, and $\hat{\epsilon}_{\phi}$ is the predicted noise. We refer readers to the Appendix for more technical details on SDS optimization.
@@ 4.1 Physics Incorporation
As mentioned in §$\S$ 3.2, we predict the dynamic behavior of generated models by rigid body simulations. While various explicit or implicit representations of 3D shapes can be chosen in a generation network, we opt for triangular meshes in our framework as they facilitate frictional contact modeling and simplify kinematics computation. Given a triangle surface mesh representation ${\mathbf{X}}(\theta)$, where $\theta$ is the implicit parameter, we integrate ${\mathbf{X}}$ into a rigid body represented by the dynamic state $\Psi(t)=[{\mathbf{T}}(t),{\mathbf{R}}(t),{\mathbf{P}}(t),{\mathbf{L}}(t)]^{T}$, where the world-space location ${\mathbf{x}}$ of any point ${\mathbf{X}}$ on the body is ${\mathbf{x}}(t)={\mathbf{R}}(t){\mathbf{X}}+{\mathbf{T}}(t)$. Assuming the 3D model is initially placed upright [^1] We define the upright pose by setting the upward axis to coincide with the upward axis from the pre-trained text-to-3D model, as the upright direction is semantics-driven and generative models inherently learn these semantics from the training data. on the ground with the bottom point touching the surface, we define standability as:
@@ 4.1.2 Stable Equilibrium
Inspired by the concept of a potential well [13], we augment our optimization objective with a robust and efficient stable equilibrium loss $\mathcal{L}_{\text{stable}}$. Specifically, for an object to be robustly standable, it needs to reside at a local minimum of potential energy—specifically, gravitational potential energy—so that if perturbed, gravity will act as a restoring force that returns the object to its original state. For a rigid body, gravitational potential energy is determined by the height of its center of mass. Thus, for any object in a stable equilibrium state, the center of mass would rise if it is slightly perturbed. This leads to our formulation of the stable equilibrium loss $\mathcal{L}_{\text{stable}}$. Let ${{\mathbf{x}}}_{\text{com}}$ denote the position of the center of mass of the underlying geometry and $H({{\mathbf{x}}})$ denote the distance of the point ${{\mathbf{x}}}$ to the ground, assuming the object’s pivot point is at $z=0$. The stable equilibrium loss is defined as:
@@ 5.1 Simulation Verification
Since we do not assume a specific baseline model in the first stage, we can vary the model used in the coarse stage to generate versatile, physically-aware 3D meshes based on different existing SDS generation models. We use MVDream [69] as another baseline model and compare the results side-by-side with ours in Fig. 5. Our method improves the geometry of the mesh and ensures stability in simulation. The quality of the texture and the main part of the shape is determined by the underlying model used in the coarse stage, while we focus on improving the physical stability in the refinement stage. More qualitative comparisons with previous methods are provided in the Appendix.
Figure 6: Ablation study of each loss term.
Ablation Study
We perform ablation studies to demonstrate the necessity of our proposed losses. It can be observed in Fig. 6(b), that without standability loss $L_{\text{stand}}$, the model fails to stand. While the model can still stand without stable equilibrium loss $L_{\text{stable}}$, as demonstrated in Fig. 6(c), it is less stable under perturbation (see next section for details). The geometry regularization loss term $L_{\text{b-lap}}$ helps smooth the geometry and avoid spiky artifacts on the surface as shown in Fig. 6(d). Additionally, we show that applying mesh regularization as a post-processing step, rather than integrating it with SDS loss in a joint optimization, can degrade text alignment as this neglects the semantics during the deformation process (see Fig. 6(e)).
Table 1: Comparison of success rate under perturbation (goose).
Perturbation Angle $\theta_{\text{max}}$ 0 0.01 0.02 0.04 0.08
w/ stability loss 1 1 0.99 0.69 0.4
w/o stability loss 1 0.97 0.71 0.62 0.23
Figure 7: Success rate of models standing under perturbation.
@@
To mimic this uncertainty in our framework, we evaluate the stability of our generated models under a small initial rotation. With a given precision $\phi_{\max}$, we rotate the generated mesh at random angles $\phi_{y},\phi_{x}\in(-\phi_{\max},\phi_{\max})$ in both the $y$ and $x$ axes, respectively (with the $z$ axis being the up direction). The mesh is then placed close to the ground and tested in a simulation to see if it can still stand. We choose 13 different values of the maximum perturbation angle $\phi_{\max}$ and perform 100 random tests with each angle on 6 of our generated models. We report the success rate in Fig. 7. We define a successful test as: after a sufficiently long time period, the maximum height of the model stays within 3% of the initial maximum height.
Due to the presence of physical constraints, our generated models can withstand small initial perturbations, while the baseline models fail to stand when placed straight up (without rotation), let alone with perturbations. Furthermore, introducing the stable equilibrium loss consistently increases the success rate of standing under different scales of perturbations, as shown in Table. 1.
@@
To evaluate the standability of the baseline method and our method, we run the rigid body simulation in Warp with simulation end time $T=2.0$ at which almost all objects have reached the steady state. We propose Time-Averaged Rotation Deviation Loss (TRD) defined as
@@
to assess the standability, as a representation of the average tilting of the upward direction $\hat{{{\mathbf{z}}}}$ (of the object) over time. We approximate the integral (Eq. 11) with discrete quadrature $\Delta t=0.02.$ Results are plotted in Fig. 9  [^2] Results are ranked by TRD scores of the baseline approach.. The mean TRD score is reduced by more than six times compared with the baseline method. As shown in Table. 2, we calculate the average CLIP score [63] of 107 generated shapes for both our proposed method and baseline. Furthermore, Elo (GPT-4o) [90] scores are presented. Both metrics illustrate that our method not only implements physics-based stability but also maintains the fidelity of the generated 3D shapes in terms of content alignment and overall shape quality. More details are provided in the Appendix.
@@ 6 Conclusion
We present Atlas3D, a physically constrained SDS-based framework that generates self-supporting 3D models from text prompts. Our framework can learn standability through a differentiable physics-based simulator and other physics-inspired loss functions. The generated 3D models can be directly imported into a physics simulator and are ready to be manufactured and deployed in the real world. Our method has wide potential for generation tasks, as it can be easily integrated into many existing pipelines and improve the physical plausibility of their generated results.
@@ Appendix A Appendix
A.1 Score Distillation Sampling
Score Distillation Sampling (SDS) has proven to be an effective method for distilling 3D models from 2D images. SDS-based methods leverage a combination of 3D representations, differentiable rendering, and pre-trained text-to-image diffusion models to optimize 3D shapes with high fidelity and realism.
The generation process in SDS methods involves optimizing the parameters $\theta$ of the 3D representation. At each iteration, the 3D geometry is differentiably rendered into a 2D image $\bm{z}=g(\theta)$. This image is then compared with the distribution of real images as modeled by the diffusion model. More specifically, the input image $\bm{z}$ is first noised with a random noise $\epsilon$ at a specified noise level $t$. The diffusion model then predicts the noise $\hat{\epsilon}_{\phi}$ with text guidance $y$, and this prediction is compared with $\epsilon$, resulting in the following loss:
$\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t,\epsilon}\left[w(t)\left\|\hat{\epsilon}_{\phi}(\bm{z}_{t},y,t)-\epsilon\right\|^{2}\right],$ (12)
where $w(t)$ is a weighting function modulating the influence of different noise levels. The expectation is taken over the noise level $t$ and noise term $\epsilon$, ensuring that the generated 3D shape aligns with the text-guided distribution of images.
The gradient with respect to the optimization parameter $\theta$ is then backpropagated as follows:
$\nabla\mathcal{L}_{\text{SDS}}(\bm{z}=g(\theta))=\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}(\bm{z}_{t},y,t)-\epsilon\right)\frac{\partial\bm{z}}{\partial\theta}\right],$ (13)
where the U-Net Jacobian term $\frac{\partial\hat{\epsilon}_{\phi}}{\partial\bm{z}_{t}}$ is omitted for efficient optimization.
For further technical details on SDS optimization, we refer readers to [58; 35; 46].
A.2 Implementation and Training Details
@@
In our experiments, we use the following default weights for the loss terms: $\{\lambda_{\text{SDS}}=1,\;\lambda_{\text{normal}}=10^{4},\;\lambda_{\text{stand}}=10^{5},\;\lambda_{\text{stable}}=10^{5},\;\lambda_{\text{b-lap}}=10^{7}\}.$ For some examples, we tune these weights within the following ranges: $\{\lambda_{\text{stand}}=10^{5}\sim 5\times 10^{5},\;\lambda_{\text{stable}}=10^{5}\sim 5\times 10^{5},\;\lambda_{\text{b-lap}}=10^{6}\sim 10^{7}\}.$ Our heuristic intuition is to keep the SDS and physical loss terms roughly on the same scale. For the regularization terms, we scale them to around $1/1000$ to $1/100$ of the SDS and physical loss terms.
A.3 Comparison with Post-processing Methods
While directly post-processing 3D generated models is a straightforward and effective approach to achieving physical stability, it may result in undesirable outcomes such as misalignment with the text prompt, as it overlooks semantics.
One simple post-processing method is to cut the mesh by a flat plane slightly higher than the lowest vertex. However, this method will fail when the projection of the center of mass lies outside the contact region, as shown in Fig. 11. Additionally, determining the cutting height is another parameter that must be manually set or optimized, and this may also degrade the overall appearance.
Figure 11: Comparison with cutting the mesh by a flat plane at height $z$.
Another post-processing method, make-it-stand [59], offers an effective way to relocate the center of mass to achieve standability. However, it assumes that the supporting surface is fixed during optimization, which can lead to distorted results due to the imperfect quality of text-to-3D generated models. We provide two examples in footnote 4. For the goose example generated by Magic3D, one leg is shorter than the other. Make-it-stand only treats one foot as the supporting surface and ignores the other due to its post-processing nature, whereas our joint-optimization pipeline enables stable standing with two legs on the ground. A similar issue also happens to the kangaroo example. More importantly, the text alignment degrades as semantics are overlooked during optimization. In contrast, our proposed joint optimization method preserves the text alignment and dynamically adjusts the center of mass as well as the supporting surface configuration.
Figure 12: Comparison with make-it-stand. [^4] Although the output mesh of make-it-stand has the same resolution as the input mesh, the ordering of vertices and surfaces are shuffled in their implementation. As a result, we are unable to re-attach the original textures onto the output mesh. We are therefore unable to run metrics like CLIP in the paper to compare text alignment. Nevertheless, it should be visually clear that our results do have better text alignment.
In Fig. 6(e) in the main text, we also apply our proposed losses in a post-processing manner. Similarly, while it is able to optimize the 3D models to make it standable, the text alignment is compromised. Overall, compared to the post-processing method, joint optimization is a more robust way which better balances text alignment and physical constraints.
A.4 Additional Results
A.4.1 Qualitative Comparison with Magic3D Baseline
@@
A.4.2 Qualitative Comparison with MVDream Baseline
@@
A.4.3 Mesh Topology Change
We provide a zoomed-in view of the local mesh topology change in Fig. 17 to demonstrate how our method optimizes the object’s geometry to make it standable.
Figure 17: Local mesh topology change
A.4.4 Quantitative Comparison with Magic3D Baseline
The results of the quantitative experiments are shown in 2. For CLIP score calculation, we specifically employ openai/clip-vit-large-patch14-336 as the check point of CLIP model. We rendered images from various angles (0-360 degree, 3 degree as the interval). For Elo (GPT-4o) benchmark, we made use of the newly released GPT-4o model instead of the one in [90] due to capacity limitation. Additionally, we reduce the number of views to 6 because of token length limitation.
A.5 Real-world Experiments
A.5.1 Robot Manipulation Results [^5] We define as upward pose for the kangaroo figure as its tail touching the ground, as a real kangaroo does.
@@ A.5 Real-world Experiments
A.5.2 User Study Results